OpenStudy (anonymous):

whats the way to solve this problem? l4x+1l=6

OpenStudy (anonymous):

You have to consider the two versions of the odulus function, the positive and negative. Firstly try the positive, 4x+1=6 x=5/4 then the negative, -4x-1=6 4x=-7 x=-7/4

OpenStudy (anonymous):

4x+1 = 6 or 4x+1 = -6 Solving the first you get x = 5/4 Solving the second you get x = -7/4 These are your solutions.

OpenStudy (anonymous):

14x=6-11 x=-5/14